The ratio of three numbers is 2:3:7. If the sum of the numbers is 360, what is the difference between the largest and the smallest numbers?
Answer by ikleyn(52750)

.

From the condition, the numbers are 2x, 3x and 7x, where x is the common measure of these numbers.


The value of x is unknown now, but can be easily foynd from the equation for the total


    2x + 3x + 7x = 360,

    12x = 360,

      x = 360/12 = 30.


Then the difference between the largest and the smallest numbers is


    7x - 2x = 5x = 5*30 = 150.     ANSWER

Solved.